OpenStudy (anonymous):

Are x=-5 and x=9 solutions to x^2-4x-45=0. please show the work. thanks

OpenStudy (anonymous):

\[x^2-4x-45=0\] \[(-5)^2-4(-5)-45=0\] \[25+20-45=0\] \[45-45=0\] \[0=0\] So, the first one looks good! :)

OpenStudy (anonymous):

\[x^2-4x-45=0\] \[(9)^2-4(9)-45=0\] \[81-36-45=0\] \[45-45=0\] \[0=0\] And the second one looks good too! :)

OpenStudy (anonymous):

Did that make sense?

OpenStudy (anonymous):

ello?

OpenStudy (anonymous):

Yes, Thank you!